12. Below are the range and standard deviation for a set of data. Use the range rule of thumb and compare it to the standard deviation listed below. Does the range rule of thumb produce an acceptable​ approximation? Suppose a researcher deems the approximation as acceptable if it has an error less than​ 15%.

Doi à https//www.mathxd.com/Student/PlayerHomeworkaspx?homeworkld-5124781868questionld-3&flushed-false... Laci Henegar Google Chrome MATH 2193 Internet Statistics Spring 2019 Homework: Section 3.3 Homework Score: 0 of 1 pt 3.3.29 Save 12 of 23 (8 complete) HW Score: 34.78%, 8 of 23 pts Question Help * Below are the range and standard deviation for a set of data. Use the range rule of thumb and compare it to the standard deviation listed below Does the range rule of thumb produce an acceptable approximation? Suppose a researcher deems the approximation as acceptable if it has an error less than 15% Range 2.83 Standard Deviation 0872 The estimated standard deviation is (Round to three doamal places as needed ) Enter your answer in the answer box and then click Check Answer remaining Clear Al Type here to search

The estimated standard deviation is___

(Round to three decimal places as​ needed.)

s this an acceptable​approximation?

A:.No​, because the error of the range rule of​ thumb's approximation is

greater than​ 15%.

B.Yes​, because the error of the range rule of​ thumb's approximation is

less than​ 15%.

C. Yes​, because the error of the range rule of​ thumb's approximation is

greater than​ 15%.

D:.No​, because the error of the range rule of​ thumb's approximation is

less than​ 15%.
Answer #1

Using range rule of thumb std deviation =range/4=2.83/4= 0.7075

Error=0.872-0.708/0.708*100

= 0.164/0.708*100

=0.01242887*100

=1.242887

<15%

ANSWERS

the estimated standard deviation is 0.708

B.Yes​, because the error of the range rule of​ thumb's approximation is

less than​ 15%.